A rectangle is 6 times as long as it is wide. The area is 150 square feet. What is the perimeter of the rectangle?

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
(6w)w = 150ft^2
w^2 = 25 (Tossing out the negative solution for unit measure)
w = 5ft and L = 30ft
P = 2*5ft + 2*30ft = 70ft

Answer by stanbon(75887)

A rectangle is 6 times as long as it is wide. The area is 150 square feet. What is the perimeter of the rectangle?
------------------------
Width:: w
Length = 6w
------
Equation:
w*6w = 150
6w^2 = 150
w^2 = 25
width = 5
-----
length = 30
------
Perimeter = 2(length + width) = 2(35) = 70 ft.
----------------
Cheers,
Stan H.
----------------